OpenStudy (anonymous):

~~~~~~~~~~~FAN AND MEDAL!~~~~~~~~~~~~~~~~ What are the x-intercept(s) of the graph of y − 3 = x2 + 4x? (−3, 0) and (−1, 0) (3, 0) and (1, 0) (−3, 0) and (1, 0) (3, 0) and (−1, 0)

OpenStudy (anonymous):

It's just a somewhat convoluted way to ask you to solve the quadratic equation. The x-intercept is where y = 0 so you're just solving \(0 - 3 = x^2 + 4x\) In other words \[x^2 - 4x + 3 = 0\] Use whatever method you like best. (I tend to prefer the quadratic formula, but this one will factor nice enough I guess)

OpenStudy (anonymous):

Oops, I copied it wrong. \(x^2 + 4x + 3\)

OpenStudy (anonymous):

lol its okay, can you show me how we would put my question into that formula?

OpenStudy (anonymous):

Ok, the quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Does that look familiar?

OpenStudy (anonymous):

Its coming back to me a bit

OpenStudy (anonymous):

The standard form of the equation is \(ax^2 + bx + c\) which is where we get those values. So for \(x^2 + 4x + 3\) it's just: \[x = \frac{-4 \pm \sqrt{4^2 - 4(1)(3)}}{2(1)}\] So just crunch that and you'll get your two solutions.

OpenStudy (anonymous):

i keep getting negative 2 @SACAPUNTAS

OpenStudy (anonymous):

@GirlByte

OpenStudy (anonymous):

\[\frac{-4 \pm \sqrt{16 - 12}}{2}\]\[\frac{-4 \pm \sqrt{4}}{2}\]\[\frac{-4 \pm 2}{2}\]\[-2 \pm 1\] So -1 and -3.

OpenStudy (anonymous):

(-3,0) (-1,0)

OpenStudy (anonymous):

Yes.